Fermi National Laboratory


Accelerator Update

Wednesday March 10
The day shift began with Operations monitoring store 3283, with Antiproton Source (Pbar) stacking, with a Recycler stash, and with MiniBooNE taking beam.

At 9:12 AM, the I- Source high voltage tripped off. It reset okay.

Operations began setting up beam for Switchyard (SY) at 9:30 AM.

The Collider Detector at Fermilab (CDF) turned off part of their detector at 11:36 AM because of problems with the colliding beam.

Operations was having a problem with getting beam through SY, but at 12:13 PM they reported seeing beam on the Meson SWICS.

At 1:30 PM, a Linac expert switched the Preaccelerator (Preacc) from the I- Source to the H- Source. The expert finished tuning the source at 3:13 PM.

A Main Injector (MI) expert began a one-hour study period at 2:30 PM.

Operation reported at 2:57 PM that the Recycler stash was gone. Experts began their investigation. (They never found a good reason for the loss.)

At 3:36 PM, the Tevatron (TeV) suffered a sector E1 quench due to a collimator that moved into the beam during an alignment.

Operations reported at 3:55 PM that the Preacc vacuum pumps kept tripping off. Operators contacted an expert.

At 4 PM operators began preparing CDF and D-zero (D0) for an access. D0 came out of access at 5 PM and CDF came out at 5:40 PM.

Operations put the TeV through a dry squeeze at 7:06 PM and then into shot setup at 10:22 PM.

Operations shut off beam for SY at 10:24 PM.

At 11:48 PM, a TeV RF station (TRF4) tripped off due to a water heater ground fault. Operators reset the station, but this required Operations to reload protons into the TeV.

Thursday March 11
The midnight shift began with the TeV, MI, and Pbar in shot setup, and with MiniBooNE taking beam.

At 12:27 AM, just after the final antiprotons were loaded, the TeV suffered a sector A1 quench. Operators began investigating the cause and contacting experts.

Pbar resumed stacking at 12:35 AM.

MI suffered a vacuum burst in its RF section at 5:23 AM. Everything reset okay.

Operations put the TeV through a dry squeeze at 8:18 AM.

At 11:20 AM, a Pbar expert began a one-hour Lithium Lens study.

Operations put the TeV into shot setup at 2:20 PM.

At 3:33 PM, the low beta magnets at sector D0 tripped off. Investigations eventually discovered that the fault was due to a failed power supply module. Coincidentally, this module had failed in a Recycler power supply last week. Power supply technicians also replaced the B0 module.

Pbar returned to stacking at 4 PM and Recycler began a study period.

Operations put the TeV into shot setup at 7:41 PM.

At 8 PM, the TeV suffered a sector E4 kicker high voltage power supply fault. This was due to a problem with low pressure on the kicker's dump resister cooling. With the help of an expert, operators got the kicker on and resumed shot setup at 8:49 PM.

Operations established store 3289 at 9:54 PM. Its initial luminosity was 47.61E30. Pbar resumed stacking.

Operation began working at establishing beam to the Meson Center (MCenter) experiment at 10:22 PM. They had beam to the experiment about forty-five minutes later.

Friday March 12
The midnight shift began with Operations monitoring store 3289, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:37 AM, a Debuncher RF station (DRF1-4) tripped off and wouldn't reset. Since Pbar could run without this station, repairs were planned for later in the morning.

A Recycler expert began setting up to stash antiprotons from Pbar at 2:11 AM.

Operations shut off SY beam to tune MI for the Recycler shots at 3:04 AM.

By 3:54 AM, the Recycler had stashed three transfers and Pbar had resumed stacking.

Operations re-established beam to SY and MCenter at 4:16 AM.

Beam problems caused CDF and D0 to put their detectors in standby at 5:42 AM. Operators tuned the beam and both experiments resumed taking data at 6:17 AM.

Today's Plans
The plan for today and this weekend is to stack and store through to Sunday night. Machines may conduct pre-shutdown studies during this time.

On Monday, March 15th, Fermilab will begin a ten-day shut to repair the TeV sector A4 vacuum leak. Of course, other machines will use the time to conduct maintenance and upgrades.

Safety system techs will conduct mandatory system tests starting at midnight on Monday. These tests will occur in many areas and run through most of the midnight shift. Actual access to any area won't be allowed until 8 AM.

Hopefully, the TeV cryo system cooldown will begin on Saturday, 3/20/04, and finish by Thursday, 3/25/04, and TeV commissioning will begin on Friday, 3/26/04.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/12/2004   email Fermilab